# Speed, Accuracy, and Complexity

### v2

@@ Abstract
This paper re-examines the validity of using response time to infer problem complexity. It revisits a canonical Wald model of optimal stopping, taking signal-to-noise ratio as a measure of problem complexity. While choice quality is monotone in problem complexity, expected stopping time is inverse $U$-shaped. Indeed decisions are fast in both very simple and very complex problems: in simple problems it is quick to understand which alternative is best, while in complex problems it would be too costly — an insight which extends to general costly information acquisition models. This non-monotonicity also underlies an ambiguous relationship between response time and ability, whereby higher ability entails slower decisions in very complex problems, but faster decisions in simple problems. Finally, this paper proposes a new method to correctly infer problem complexity based on the finding that choices react more to changes in incentives in more complex problems.
@@ 1.  Introduction
It is almost definitional that more complex decision problems entail more mistakes. Individuals make more mistakes and choose dominated alternatives in more complex environments, both in experimental settings (e.g. Rabin and Weizsäcker, 2009; Martínez-Marquina et al., 2019; Banovetz and Oprea, 2023; Enke and Graeber, 2023; Puri, 2023) and in real-world decisions with significant financial implications.111 Examples include choosing health insurance plans (Sinaiko and Hirth, 2011; Bhargava et al., 2017), pension plans (Choi et al., 2011), and mortgages (Keys et al., 2016; Agarwal et al., 2016). This has important implications, since deploying theoretically optimal but complex incentive structures (tax schedules, allocation mechanisms, pricing schemes, etc) may lead to unreliable predictions and suboptimal outcomes — an observation at the root of the ongoing focus on simple mechanisms, both in theoretical and in empirical work.222 For recent theoretical work concerned with simplicity in mechanism design see Li (2017), Börgers and Li (2019), and Pycia and Troyan (2023); an example of empirical research in the same domain is Rees-Jones and Taubinsky (2020). If problem complexity is taken to be a leading explanation of suboptimal decision-making, understanding which problems are more complex becomes crucial for simplifying them.333 Or to obfuscate them further: a growing literature discusses how complexity can be leveraged and manipulated to the advantage of principals — cf. Ellison and Wolitzky (2012), Jehiel (2011), Spiegler (2016), and Li and Dworczak (2022).
@@
Response time is often used as an indicator of problem complexity in various settings, including lottery choice (Wilcox, 1993), loans (Kalaycı and Serra-Garcia, 2016), and portfolio selection (Carvalho and Silverman, 2017).444 Many of these studies show a correlation between response time and problem complexity, rather than assuming it. Other recent examples are Frydman and Nunnari (2023), who use decision time together other behavioural patterns in support of specific coordination settings being more complex, and Horenstein and Grabiszewski (2022), who use it to rank extensive-form games in terms of complexity. More examples are surveyed in Spiliopoulos and Ortmann (2018) and Gill and Prowse (2023). Going beyond economics, a large literature in psychology and neuroscience has also used response time as a proxy for problem complexity (Wright and Ayton, 1988; Ratcliff and McKoon, 2008; Shadlen and Shohamy, 2016; Bogacz et al., 2010). The underlying assumption is that in more complex problems, making similarly good choices requires greater cognitive effort, and response time can serve as a proxy for cognitive effort (Caplin et al., 2020). If decision-makers are able to make more accurate choices while expending less cognitive effort, as indicated by faster responses, then the decision problem is considered simpler. Therefore, a decision problem is deemed simpler if individuals make fewer mistakes (more accurate choices) and do so with less cognitive effort (faster choices).
@@
The reliance on response time as a proxy for problem complexity is based on the observation that, in many cases, faster choices tend to be more accurate. For example, research in cognitive sciences has indicated that perceptual problems that are harder to discern require greater cognitive effort, leading to slower and less accurate decisions (e.g. Forstmann et al., 2016). Additionally, studies have found that decision-makers take longer when they are closest to being indifferent, indicating the difficulty in distinguishing the best alternative (Mosteller and Nogee, 1951; Krajbich et al., 2012; Clithero, 2018; Alós-Ferrer and Garagnani, 2022). This intuition aligns with sequential sampling models (Fudenberg et al., 2018; Gonçalves, 2023; Alós-Ferrer et al., 2021), where decisions take longer when individuals are closer to being indifferent, as the value of information is highest in such cases.555 See also Drugowitsch et al. (2012). In short, the justification behind the use of response times as a proxy for problem complexity is straightforward: a problem is considered easier if choices are both faster and more accurate, with faster choices often implying greater accuracy.
@@
Whilst response time may be a compelling indicator of problem complexity, it may not always be a reliable measure when used on its own. Contributing to an long-standing discussion (Hogarth, 1975; Ford, 1975; Onken et al., 1985; Wright and Ayton, 1988), a more recent literature documents a more ambiguous relationship between speed and accuracy, where in some cases faster decisions are better whereas in others slower decisions are more accurate (e.g. Achtziger and Alós-Ferrer, 2013; Moritz et al., 2014; Caplin and Martin, 2016). Introspection also cautions against positing a monotone relationship between speed, accuracy, and complexity: if problem complexity demands greater cognitive effort and longer decision time, then in excessively complex problems thinking for a long time may not be worthwhile.
@@
Our baseline setup is the most canonical sequential sampling model (Wald, 1947; Dvoretzky et al., 1953), where a decision-maker faces a binary choice problem with a constant flow cost per unit of time and can obtain information about a binary event or parameter by observing a Brownian motion with drift. We take the signal-to-noise ratio, i.e. the drift intensity relative to the instantaneous volatility, as a measure of problem complexity, as is common in existing literature (e.g. Callander, 2011; Fehr and Rangel, 2011). The nature of the exercise is then to examine how speed (expected stopping time) and accuracy (probability of choosing the optimal alternative) depend on problem complexity.
@@
Several papers have examined the relation between response time and ability. For instance, it is often found that slower responses are indicative of higher ability or sophistication (e.g. Rubinstein, 2016; Agranov et al., 2015; Recalde et al., 2018; Esteban-Casanelles and Gonçalves, 2022). However, other studies suggest that, similarly to accuracy, this relationship is ambiguous, with greater ability being associated with faster response times in simple problems (Dodonova and Dodonov, 2013; Goldhammer et al., 2014; Kyllonen and Zu, 2016). Anecdotal evidence again cautions against generic claims for monotone association: for instance, returning exams earlier typically indicates that a student will score significantly higher or significantly lower than their peers.
@@
We show that this ambiguity is a natural consequence of the non-monotone relation between speed and accuracy. For this we extend our model to allow for the decision-maker to control instantaneous effort, which entails a cost per unit of time. In our model, higher effort scales up the signal-to-noise ratio,777 Our setup is akin to Moscarini and Smith (2001), but with no discounting, to keep the model as close as possible to our baseline. and the decision-maker will optimally choose a constant level of effort. This can then be interpreted as if the decision-maker faces an idiosyncratic problem with complexity level that depends on their cost of effort. We contrast the speed and accuracy of high-ability and low-ability decision-makers, where higher ability captures a lower cost of effort and show that high ability individuals always choose better, but do so faster than lower ability individuals in simple problems, and slower in more complex ones. Furthermore, this also cautions against pooling data on multiple individuals, since then response time will not in general be single-peaked.
@@
We extend our results to more general models of costly information acquisition. Going to a more general setup, we consider general costly information acquisition models (e.g. Caplin and Dean, 2015), where the decision-maker can acquire a Blackwell experiment to learn about an unknown state paying a cost that is increasing with respect to the Blackwell order. We assume that (1) problem complexity corresponds to a parameter that scales up the cost of information,888 As is standard in costly information acquisition or rational inattention models with entropy costs (Matějka and McKay, 2015, e.g.,). and (2) that expected stopping time is proportional to the total cost of information. The first assumption is motivated by the fact that, in the case of the Wald problem, the square of noise-to-signal ratio corresponds to such parameter (Morris and Strack, 2019), whereas the second assumption not only exactly corresponds to our baseline setup, but is also consistent with evidence from Caplin et al. (2020). We then show that, in this general setup, expected stopping time is then non-monotone in problem complexity, and provide conditions for it to be single-peaked.999 While we only consider models in which the cost information is continuous and convex, recent literature has also indicated that in binary decision problem with one-sided Poisson process, the expected decision time could be non-monotone in the arrival rate (Bobtcheff and Levy, 2017; Halac et al., 2016). This non-monotonicity, however, does seem to depend on modelling details — see Bobtcheff and Levy (2017, Proposition 1).
@@ 2.  Setup
A decision-maker is choosing from a set $A=\{a,b\}$, with generic element $\alpha\in A$. Let $(\Omega,\mathcal{F},\mathbb{P})$ denote a probability space, where $\omega\in\Omega$ denotes the state of the world, $\mathcal{F}$ a Borel $\sigma$-algebra, and $\mathbb{P}$ a probability measure. Payoffs to each alternative depend on a parameter or event $\theta\in\Theta:=\{a,b\}$, and are given by $u:A\times\Theta\to\mathbb{R}$. We assume there are not weakly dominant alternatives so that the choice problem is non-trivial and, without loss of generality, $\theta$ specifies which alternative is best, i.e. $u(\theta,\theta)>u(\theta^{\prime},\theta)$, $\theta\neq\theta^{\prime}$.
@@
The model is typically interpreted to be a stylised representation of the process of deliberation or reasoning in decision-making, with the state of the process $X_{t}$ meant to represent accumulated evidence in favour of or against any of the alternatives. While in psychology and neuroscience it became a benchmark in modelling the neural basis of decision-making,131313 For instance, it has been the main modelling paradigm for the recalling past experiences and sampling from memory (Ratcliff, 1978; Shadlen and Shohamy, 2016), the sensory perception of physical characteristics (Gold and Shadlen, 2001; Ratcliff et al., 2016), and brain activation in decision-making (Bogacz et al., 2010; Shushruth et al., 2022). Recent variations extend the model to interpret eye gase, e.g. in purchasing decisions (Krajbich et al., 2012). in economics it has been used to model a wide range of situations and behavioural patterns, including randomness in consumer choice (Clithero, 2018; Alós-Ferrer et al., 2021), the relation between misperception and small-stakes risk aversion (Frydman and Jin, 2022; Khaw et al., 2021), consumer’s search for product information (Branco et al., 2012), choosing actions in games (Schotter and Trevino, 2021), information gathering and deliberation in groups (Reshidi et al., 2024), and brand recognition (Alós-Ferrer, 2018).
@@
As the signal-to-noise ratio $\mu/\sigma$ captures the rate of information flow per unit of time, we will take this as a measure of problem complexity — with higher values of $\mu/\sigma$ denoting easier problems. This is a common interpretation in economics, cognitive sciences, and psychology (e.g. Callander, 2011; Fehr and Rangel, 2011; Forstmann et al., 2016). In essence, it captures the intuition that simple problems (high $\mu/\sigma$) are those in which a decision-maker quickly realizes which alternative or course of action is best, whereas in complex problems (low $\mu/\sigma$) it would take longer to be as confident. A decision problem may be more complex owing to the intrinsic properties of the items involved — e.g. deciding between savings products is arguably easier and faster than understanding which structured financial product is best for oneself; solving easy exam questions is faster than harder ones. Alternatively, how complex a problem is may have to do with how information is presented or how easy it to obtain information, making it easier or harder to compare across alternatives.
@@ 3.1.  Characterising Optimal Stopping
Let us introduce the following auxiliary payoff functions: $u_{a}(\alpha,p):=1_{\alpha=a}\cdot(p-\tilde{p})$ and $u_{b}(\alpha,p)=1_{\alpha=b}(\tilde{p}-p)$, where $\tilde{p}$ denotes the belief that renders the decision-maker indifferent between both alternatives, i.e. $\tilde{p}:u(a,\tilde{p})=u(b,\tilde{p})$. The interpretation of these payoff functions is simple: for $\alpha\in A$, $u_{\alpha}$ is such that $\alpha$ is the risky alternative and the other is the safe one, delivering a payoff normalised to zero. This then implies that the relation between beliefs and preference is retained: $u(a,p)\geq u(b,p)\Longleftrightarrow u_{a}(a,p)\geq u_{a}(b,p)\Longleftrightarrow u_{b}(a,p)\geq u_{b}(b,p)\Longleftrightarrow p\geq\tilde{p}$.
@@ 3.2.  Speed and Accuracy
Theorem 1 also implies that easier problems translate into more accurate choices, which, as the proof shows, is due to wider stopping thresholds for beliefs. While wider thresholds entail longer expected stopping times, a higher signal-to-noise ratio entail faster choices. The former effect dominates for more difficult problems, when the signal-to-noise ratio is low, whereas the latter dominates when simpler problems are made simpler. This result is consistent with existing evidence that, in simple situations, faster choices are more accurate, but in more complicated ones, the opposite is true (e.g. Onken et al., 1985; Achtziger and Alós-Ferrer, 2013).
We want to highlight two common variants of the model that will not affect this observation. In many instances, the drift is proportional to the payoff difference (e.g. Fehr and Rangel, 2011). In our model, this would be observationally equivalent to having a drift rate of $\mu\;\delta$ rather than just $\mu$. Theorem 1 would hold just the same in this case. A second variant of the model would feature exponential discounting instead of a flow cost of time; in Appendix B we show that our results are robust: the relation between accuracy and expected stopping time remains non-monotone.
One important implicit assumption in our analysis is that the decision-maker can to some extent recognise the complexity of the problem they face. In this, we differ from previous literature, e.g. Fudenberg et al. (2018), who preclude this possibility. While it may often be the case that one learns about how complicated a particular problem is while trying to solve it, there may be many instances in which there are apparent differences in complexity across problems, and the decision-maker can realise this. In such cases, without further assumptions, one should not in general expect the speed of decision to indicate response accuracy or problem complexity.

### v3

@@ Abstract
This paper re-examines the use of response time to infer problem complexity. It revisits a canonical Wald model of optimal stopping, taking signal-to-noise ratio as a measure of problem complexity. While choice quality is monotone in problem complexity, expected stopping time is inverse U-shaped. Indeed, decisions are fast in both very simple and very complex problems: in simple problems, it is quick to understand which alternative is best, while in complex problems it would be too costly — an insight which extends to general costly information acquisition models. This non-monotonicity also underlies an ambiguous relationship between response time and ability, whereby higher ability entails slower decisions in very complex problems, but faster decisions in simple problems. Finally, this paper proposes a new method to correctly infer problem complexity based on the finding that distorting incentives in favour of an alternative has a greater effect on choices in more complex problems.
@@ 1.  Introduction
It is almost definitional that more complex decision problems entail more mistakes. Individuals make more mistakes and choose dominated alternatives in more complex environments, both in experimental settings (e.g., Rabin and Weizsäcker, 2009; Martínez-Marquina et al., 2019; Banovetz and Oprea, 2023; Enke and Graeber, 2023; Puri, 2023) and in real-world decisions with significant financial implications.111 Examples include choosing health insurance plans (Sinaiko and Hirth, 2011; Bhargava et al., 2017), pension plans (Choi et al., 2011), and mortgages (Keys et al., 2016; Agarwal et al., 2016). This has important implications, since deploying theoretically optimal but complex incentive structures (tax schedules, allocation mechanisms, pricing schemes, etc) may lead to unreliable predictions and suboptimal outcomes — an observation at the root of the ongoing focus on simple mechanisms, both in theoretical and in empirical work.222 For recent theoretical work concerned with simplicity in mechanism design see Li (2017), Börgers and Li (2019), and Pycia and Troyan (2023); an example of empirical research in the same domain is Rees-Jones and Taubinsky (2020). If problem complexity is taken to be a leading explanation of suboptimal decision-making, understanding which problems are more complex becomes crucial for simplifying them.333 Or to obfuscate them further: a growing literature discusses how complexity can be leveraged and manipulated to the advantage of principals — cf. Ellison and Wolitzky (2012), Jehiel (2011), Spiegler (2016), and Li and Dworczak (2022).
@@
Response time is often used as an indicator of problem complexity in various settings, including lottery choice (Wilcox, 1993), loans (Kalaycı and Serra-Garcia, 2016), and portfolio selection (Carvalho and Silverman, 2017).444 These studies show — rather than assuming it — a correlation between response time and problem complexity. Other recent examples are Frydman and Nunnari (2023), who use decision time together other behavioural patterns in support of specific coordination settings being more complex, and Horenstein and Grabiszewski (2022), who use it to rank extensive-form games in terms of complexity. More examples are surveyed in Spiliopoulos and Ortmann (2018) and Gill and Prowse (2023). Going beyond economics, a large literature in psychology and neuroscience has also used response time as a proxy for problem complexity (Wright and Ayton, 1988; Ratcliff and McKoon, 2008; Shadlen and Shohamy, 2016; Bogacz et al., 2010). The underlying assumption is that in more complex problems, making similarly good choices requires greater cognitive effort, and response time can serve as a proxy for cognitive effort (Caplin et al., 2020). If decision-makers are able to make more accurate choices while expending less cognitive effort, as indicated by faster responses, then the decision problem is considered simpler. Therefore, a decision problem is deemed simpler if individuals make fewer mistakes (more accurate choices) and do so with less cognitive effort (faster choices).
@@
The reliance on response time as a proxy for problem complexity is based on the observation that, in many cases, faster choices tend to be more accurate. For example, research in cognitive sciences has indicated that perceptual problems that are harder to discern require greater cognitive effort, leading to slower and less accurate decisions (e.g., Forstmann et al., 2016). Additionally, studies have found that decision-makers take longer when they are closest to being indifferent, indicating the difficulty in distinguishing the best alternative (Mosteller and Nogee, 1951; Krajbich et al., 2012; Clithero, 2018; Alós-Ferrer and Garagnani, 2022). This intuition aligns with sequential sampling models (Fudenberg et al., 2018; Gonçalves, 2023; Alós-Ferrer et al., 2021), where decisions take longer when individuals are closer to being indifferent, as the value of information is highest in such cases.555 See also Drugowitsch et al. (2012). In short, the justification behind the use of response times as a proxy for problem complexity is straightforward: a problem is considered easier if choices are both faster and more accurate, with faster choices often implying greater accuracy.
@@
Whilst response time may be a compelling indicator of problem complexity, it may not always be a reliable measure when used on its own. Contributing to an long-standing discussion (Hogarth, 1975; Ford, 1975; Onken et al., 1985; Wright and Ayton, 1988), a more recent literature documents a more ambiguous relationship between speed and accuracy, where in some cases faster decisions are better whereas in others slower decisions are more accurate (e.g., Achtziger and Alós-Ferrer, 2013; Moritz et al., 2014; Caplin and Martin, 2016). Introspection also cautions against positing a monotone relationship between speed, accuracy, and complexity: if problem complexity demands greater cognitive effort and longer decision time, then in excessively complex problems thinking for a long time may not be worthwhile.
@@
Our baseline setup is the most canonical sequential sampling model (Wald, 1947; Dvoretzky et al., 1953), where a decision-maker faces a binary choice problem with a constant flow cost per unit of time and can obtain information about a binary event or parameter by observing a Brownian motion with drift. We take the signal-to-noise ratio, i.e., the drift intensity relative to the instantaneous volatility, as a measure of problem complexity, as is common in existing literature (e.g., Callander, 2011; Fehr and Rangel, 2011). The nature of the exercise is then to examine how speed (expected stopping time) and accuracy (probability of choosing the optimal alternative) depend on problem complexity.
@@
Several papers have examined the relation between response time and ability. For instance, it is often found that slower responses are indicative of higher ability or sophistication (e.g., Rubinstein, 2016; Agranov et al., 2015; Recalde et al., 2018; Esteban-Casanelles and Gonçalves, 2022). However, other studies suggest that, similarly to accuracy, this relationship is ambiguous, with greater ability being associated with faster response times in simple problems (Dodonova and Dodonov, 2013; Goldhammer et al., 2014; Kyllonen and Zu, 2016). Anecdotal evidence again cautions against generic claims for monotone association: for instance, returning exams earlier typically indicates that a student will score significantly higher or significantly lower than their peers.
@@
We show that this ambiguity is a natural consequence of the non-monotone relation between speed and accuracy. For this we extend our model to allow for the decision-maker to control instantaneous effort, which entails a cost per unit of time. In our model, higher effort scales up the signal-to-noise ratio,777 Our setup is akin to Moscarini and Smith (2001), but with no discounting, to keep the model as close as possible to our baseline. and the decision-maker will optimally choose a constant level of effort. This can then be interpreted as if the decision-maker faces an idiosyncratic problem with complexity level that depends on their cost of effort. We contrast the speed and accuracy of high-ability and low-ability decision-makers, where higher ability captures a lower cost of effort, and show that high ability individuals always choose better, but do so faster than lower ability individuals in simple problems, and slower in more complex ones. This also cautions against pooling data on multiple individuals, since then response time will not in general be single-peaked.
@@
We extend our results to more general models of costly information acquisition. Going to a more general setup, we consider general costly information acquisition models (e.g., Caplin and Dean, 2015), where the decision-maker can acquire a Blackwell experiment to learn about an unknown state paying a cost that is increasing with respect to the Blackwell order. We assume that (1) problem complexity corresponds to a parameter that scales up the cost of information,888 As is standard in costly information acquisition or rational inattention models with entropy costs (Matějka and McKay, 2015, e.g.,). and (2) that expected stopping time is proportional to the total cost of information. The first assumption is motivated by the fact that, in the case of the Wald problem, the square of noise-to-signal ratio corresponds to such parameter (Morris and Strack, 2019), whereas the second assumption not only exactly corresponds to our baseline setup, but is also consistent with evidence from Caplin et al. (2020). We then show that, in this general setup, expected stopping time is then non-monotone in problem complexity, and provide conditions for it to be single-peaked.999 Our results give breadth to an existing literature pointing to related findings. Steiner et al. (2017) study a dynamic rational inattention model with capacity constraints and provide numerical simulations indicating the stopping rate is non-monotone in the capacity. In binary decision problems with one-sided Poisson process, depending on the setup details (see Bobtcheff and Levy, 2017, Proposition 1), expected decision time could be non-monotone in the arrival rate.
@@ 2.  Setup
A decision-maker is choosing from a set $A=\{a,b\}$, with generic element $\alpha\in A$. Let $(\Omega,\mathcal{F},\mathbb{P})$ denote a probability space, where $\omega\in\Omega$ denotes the state of the world, $\mathcal{F}$ a Borel $\sigma$-algebra, and $\mathbb{P}$ a probability measure. Payoffs to each alternative depend on a parameter or event $\theta\in\Theta:=\{a,b\}$, and are given by $u:A\times\Theta\to\mathbb{R}$. We assume there are not weakly dominant alternatives so that the choice problem is non-trivial and, without loss of generality, $\theta$ specifies which alternative is best, i.e., $u(\theta,\theta)>u(\theta^{\prime},\theta)$, $\theta\neq\theta^{\prime}$.
@@
The model is typically interpreted to be a stylised representation of the process of deliberation or reasoning in decision-making, with the state of the process $X_{t}$ meant to represent accumulated evidence in favour of or against any of the alternatives. While in psychology and neuroscience it became a benchmark in modelling the neural basis of decision-making,131313 For instance, it has been the main modelling paradigm for the recalling past experiences and sampling from memory (Ratcliff, 1978; Shadlen and Shohamy, 2016), the sensory perception of physical characteristics (Gold and Shadlen, 2001; Ratcliff et al., 2016), and brain activation in decision-making (Bogacz et al., 2010; Shushruth et al., 2022). Recent variations extend the model to interpret eye gase, e.g., in purchasing decisions (Krajbich et al., 2012). in economics it has been used to model a wide range of situations and behavioural patterns, including randomness in consumer choice (Clithero, 2018; Alós-Ferrer et al., 2021), the relation between misperception and small-stakes risk aversion (Frydman and Jin, 2022; Khaw et al., 2021), consumer’s search for product information (Branco et al., 2012), choosing actions in games (Schotter and Trevino, 2021), information gathering and deliberation in groups (Reshidi et al., 2024), and brand recognition (Alós-Ferrer, 2018).
@@
As the signal-to-noise ratio $\mu/\sigma$ captures the rate of information flow per unit of time, we will take this as a measure of problem complexity — with higher values of $\mu/\sigma$ denoting easier problems. This is a common interpretation in economics, cognitive sciences, and psychology (e.g., Callander, 2011; Fehr and Rangel, 2011; Forstmann et al., 2016). In essence, it captures the intuition that simple problems (high $\mu/\sigma$) are those in which a decision-maker quickly realizes which alternative or course of action is best, whereas in complex problems (low $\mu/\sigma$) it would take longer to be as confident. A decision problem may be more complex owing to the intrinsic properties of the items involved — e.g., deciding between savings products is arguably easier and faster than understanding which structured financial product is best for oneself; solving easy exam questions is faster than harder ones. Alternatively, how complex a problem is may have to do with how information is presented or how easy it to obtain information, making it easier or harder to compare across alternatives.
@@ 3.1.  Characterising Optimal Stopping
Let us introduce the following auxiliary payoff functions: $u_{a}(\alpha,p):=1_{\alpha=a}\cdot(p-\tilde{p})$ and $u_{b}(\alpha,p)=1_{\alpha=b}(\tilde{p}-p)$, where $\tilde{p}$ denotes the belief that renders the decision-maker indifferent between both alternatives, i.e., $\tilde{p}:u(a,\tilde{p})=u(b,\tilde{p})$. The interpretation of these payoff functions is simple: for $\alpha\in A$, $u_{\alpha}$ is such that $\alpha$ is the risky alternative and the other is the safe one, delivering a payoff normalised to zero. This then implies that the relation between beliefs and preference is retained: $u(a,p)\geq u(b,p)\Longleftrightarrow u_{a}(a,p)\geq u_{a}(b,p)\Longleftrightarrow u_{b}(a,p)\geq u_{b}(b,p)\Longleftrightarrow p\geq\tilde{p}$.
@@ 3.2.  Speed and Accuracy
Theorem 1 also implies that easier problems translate into more accurate choices, which, as the proof shows, is due to wider stopping thresholds for beliefs. While wider thresholds entail longer expected stopping times, a higher signal-to-noise ratio entail faster choices. The former effect dominates for more difficult problems, when the signal-to-noise ratio is low, whereas the latter dominates when simpler problems are made simpler. This result is consistent with existing evidence that, in simple situations, faster choices are more accurate, but in more complicated ones, the opposite is true (e.g., Onken et al., 1985; Achtziger and Alós-Ferrer, 2013).
We want to highlight two common variants of the model that will not affect this observation. In many instances, the drift is proportional to the payoff difference (e.g., Fehr and Rangel, 2011). In our model, this would be observationally equivalent to having a drift rate of $\mu\;\delta$ rather than just $\mu$. Theorem 1 would hold just the same in this case. A second variant of the model would feature exponential discounting instead of a flow cost of time; in Appendix B we show that our results are robust: the relation between accuracy and expected stopping time remains non-monotone. Beyond this, our results extend to other flow costs, e.g., allowing the ones entailing ex-ante mutual information costs.
One important implicit assumption in our analysis is that the decision-maker can to some extent recognise the complexity of the problem they face. In this, we differ from previous literature, e.g., Fudenberg et al. (2018), who preclude this possibility. While it may often be the case that one learns about how complicated a particular problem is while trying to solve it, there may be many instances in which there are apparent differences in complexity across problems, and the decision-maker can realise this. In such cases, without further assumptions, one should not in general expect the speed of decision to indicate response accuracy or problem complexity.